OpenStudy (cryleeb):

WILL FAN AND MEDAL Consider the following pair of equations: y = –2x + 8 y = x – 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form. (5 points)

OpenStudy (jingle535):

The y's equal each other so the equations equal eachother -2x+8=x-1 Addition Property -2x+9=x Addition Property 9=3x Division Property 3=x -------------- 3=x, substituting the 3 in for the x in the equation y=3-1 y=2

OpenStudy (jingle535):

Sorry I must go now good luck :)

OpenStudy (cryleeb):

thankyou!